Which of the following vectors is perpendicular to the vector 4.0 - 6.0 +2.0?
Answer #1

We have vector X = 4i-6j+2k . Let's say vector A = 21 - 3j + 10k is perpendicular to the vector X .

If vector X and \overrightarrow{A} are perpendicular to each other then their dot product is zero:

子了=0

:: (4 - 6j + 2k). (2i - 3 + 10k) = 0

..8+18+ 20 = 0

1:46 +0

Therefore, vector À = 2 - 3j + 10k is not perpendicular to vector X = 4i-6j+2k .

We have vector X = 4i-6j+2k . Let's say vector В = 3i — 3ј+ Зk is perpendicular to the vector X .

If vector X and B are perpendicular to each other then their dot product is zero:

B 0

:: (4 - 6j + 2k) (31 - 3; +3k) = 0

.. 12-18 +6=0

. 36 ¥0

Therefore, vector В = 3i — 3ј+ Зk is not perpendicular to vector X = 4i-6j+2k .

We have vector X = 4i-6j+2k . Let's say vector & = lî - 2 + 1h is perpendicular to the vector X .

If vector X and \overrightarrow{C} are perpendicular to each other then their dot product is zero:

\therefore \overrightarrow{X}\cdot \overrightarrow{C}=0

\therefore (4\widehat{i}-6\widehat{j}+2\widehat{k})\cdot (1\widehat{i}-2\widehat{j}+1\widehat{k})=0

..4-12-2=0

. 18 ¥0

Therefore, vector {\color{Red} \overrightarrow{C}=1\widehat{i}-2\widehat{j}+1\widehat{k}} is not perpendicular to vector X = 4i-6j+2k .

We have vector X = 4i-6j+2k . Let's say vector \overrightarrow{D}=1\widehat{i}+1\widehat{j}+1\widehat{k} is perpendicular to the vector X .

If vector X and \overrightarrow{D} are perpendicular to each other then their dot product is zero:

\therefore \overrightarrow{X}\cdot \overrightarrow{D}=0

\therefore (4\widehat{i}-6\widehat{j}+2\widehat{k})\cdot (1\widehat{i}+1\widehat{j}+1\widehat{k})=0

.. 4-6+2=0

{\color{DarkGreen} \therefore 0= 0}

Therefore, vector Ď= li +1j + 1k is perpendicular to vector X = 4i-6j+2k .

We have vector X = 4i-6j+2k . Let's say vector \overrightarrow{E}=2\widehat{i}+1\widehat{j}+0\widehat{k} is perpendicular to the vector X .

If vector X and \overrightarrow{E} are perpendicular to each other then their dot product is zero:

\therefore \overrightarrow{X}\cdot \overrightarrow{E}=0

\therefore (4\widehat{i}-6\widehat{j}+2\widehat{k})\cdot (2\widehat{i}+1\widehat{j}+0\widehat{k})=0

\therefore 8-6+0=0

{\color{Red} \therefore 2\neq 0}

Therefore, vector E = 2 +1j+ok is not perpendicular to vector X = 4i-6j+2k .